# Generalized Random Forests using Fixed-Point Trees

### v3

@@ C.1 Honest subsampling
$\alpha_{i}(x)\coloneqq\frac{1}{B}\sum^{B}_{b=1}\alpha_{bi}(x),\quad\text{for}\quad\alpha_{bi}(x)\coloneqq\frac{\mathds{1}(\{X_{i}\in L_{b}(x)\})}{|L_{b}(x)|},\qquad i=1,\ldots,n,$ (36)
@@
$\mathbb{E}[\alpha_{bi}(x)\,|\,O_{i},X_{i}]=\mathbb{E}[\alpha_{bi}(x)\,|\,X_{i}].$
@@
    $\mathbb{E}[\alpha_{bi}(x)\,|\,O_{i},X_{i}]=\mathbb{E}[\alpha_{bi}(x)\,|\,X_{i}]=0,\qquad\text{for all}\quad i\notin\mathcal{J}^{(b)}_{2}.$
@@
    $\mathbb{E}[\alpha_{bi}(x)\,|\,O_{i},X_{i}]=\mathbb{E}[\alpha_{bi}(x)\,|\,X_{i}],\qquad\text{for all}\quad i\in\mathcal{J}^{(b)}_{2}.$
@@ C.4 Simulation Details
  • Implementation details. The original gradient-based implementation of GRF is available in the R package grf (Tibshirani et al., 2024). We implement the FPT algorithm in a fork of grf at https://github.com/dfleis/grf. The functions lm_forest and multi_arm_causal_forest provide an easy to use interface for VCM and HTE estimation, respectively, with the choice of GRF-FPT or GRF-grad controlled via the method argument. Code and data for reproducing all experiments and figures are available at https://github.com/dfleis/grf-experiments.
  • Implementation details. The original gradient-based implementation of GRF is available in the R package grf (Tibshirani et al., 2024). We implement the FPT algorithm in a fork of grf. The functions lm_forest and multi_arm_causal_forest provide an easy to use interface for VCM and HTE estimation, respectively, with the choice of GRF-FPT or GRF-grad controlled via the method argument.
  • Data-generating settings. The different setting for the target effects $\theta_{k}^{*}(x)$ include a sparse linear setting, a sparse logistic setting with interaction, a dense logistic setting, and a random function generator setting. Tables 2 and 3 provide the details of each regime for VCM and HTE experiments, respectively, for the data-generating model (25). These tables also summarize the different settings used to generate the $K$-dimensional regressors $W_{i}=(W_{i,1},\ldots,W_{i,K})^{\top}$. For VCM experiments, $W_{i,k}\sim\mathcal{N}(0,1)$ for all $k=1,\ldots,K$. For HTE experiments, $W_{i}|X_{i}=x\sim\text{Multinomial}(1,(\pi_{1}(x),\ldots,\pi_{K}(x)))$, where $\pi_{k}(x)$ denotes the underlying probability the sample is observed as having treatment level $k\in\{1,\ldots,K\}$.
@@
    $K$ 256
    $n$ 10,000
    $\dim(X)$ 5
    nTrees 100; 250; 500; 1000
    Table 1: Parameter values for VCM and HTE experiments under the data-generating model (25). Target/regressor dimension $K$, number of observations $n$, dimension of the auxiliary variables $\dim(X)$, and number of trees nTrees.
@@ 4 $\theta^{*}_{k}(x)=\text{RFG}(x)$ $\mathcal{N}(0,1)$
    Table 2: Settings for the underlying effects $\theta^{*}_{k}(\cdot)$ and the regressors $W_{i,k}$ for VCM experiments described in Section 6. The function $\varsigma(u)\coloneqq 1+(1+e^{-20(u-1/3)})^{-1}$ is a logistic-type function used for the simulations in (Athey et al., 2019). The random function generator $\text{RFG}(x)$ is described in Appendix C.4.
@@
    Table 3: Settings for the underlying treatment effects $\theta^{*}_{k}(\cdot)$ and treatment probabilities $\pi_{k}(x)$ for HTE experiments described in Section 6. The function $\varsigma(u)\coloneqq 1+(1+e^{-20(u-1/3)})^{-1}$ is a logistic-type function used for the simulations in (Athey et al., 2019). The random function generator $\text{RFG}(x)$ is described in Appendix C.4.
@@
  • Random function generator $\text{RFG}(x)$. For the simulations in Section 6, the effect functions $\theta^{*}_{k}(x)=\text{RFG}(x)$ follow the “random function generator” model by Friedman (2001). The idea is to measure the performance of the estimator under a variety of randomly generated targets. Each $\theta^{*}_{k}(\cdot)$ is randomly generated as a linear combination of functions $\{g_{l}(\cdot)\}_{1}^{20}$ of the form
    $\theta^{*}_{k}(x)=\sum_{l=1}^{20}a_{l}g_{l}(z_{l}),$
@@
    where the coefficients $\{a_{l}\}_{1}^{20}$ are randomly generated from a uniform distribution $a_{l}\sim\mathcal{U}([-1,1])$. Each $g_{l}(z_{l})$ is a function of a randomly selected $p_{l}$-size subset of the $p$-dimensional variable $x$, where the size of each subset $p_{l}$ is randomly chosen by $p_{l}=\min(\left\lfloor 2.5+r\right\rfloor,p)$, and $r$ is generated from an exponential distribution $r\sim\mathrm{Exp}(0.5)$ with mean $2$. Each $z_{l}$ is defined as
    $z_{l}=\{x_{W_{l}(j)}\}_{j=1}^{p_{l}},$
    where each $W_{l}$ is an independent permutation of the integers $\{1,\ldots,p\}$. Each function $g_{l}(z_{l})$ is a $p_{l}$-dimensional Gaussian function:
    $g_{l}(z_{l})=\exp\left[-\frac{1}{2}(z_{l}-\mu_{l})^{T}\mathbf{V}_{l}(z_{l}-\mu_{l})\right],$
    where each of the mean vectors $\{\mu_{l}\}_{1}^{20}$ is randomly generated from the same distribution as that of the input variables $x$. The $p_{l}\times p_{l}$ covariance matrix $\mathbf{V}_{l}$ is also randomly generated according to
    $\mathbf{V}_{l}=\mathbf{U}_{l}\mathbf{D}_{l}\mathbf{U}_{l}^{T},$
@@
    where $\mathbf{U}_{l}$ is a random orthonormal matrix and $\mathbf{D}_{l}=\text{diag}\{d_{1l}...d_{p_{l}l}\}$ where diagonal entries $d_{jl}$ are generated from a uniform distribution according to $\sqrt{d_{jl}}\sim\mathcal{U}([0.1,2.0])$.
  • Single tree timings. Subsampling and honest sample splitting was disabled for single tree timing experiments via the arguments sample.fraction = 1 and honesty = F for both VCM and HTE simulations. The focus of these experiments is on the time to grow a single tree under the GRF-FPT and GRF-grad methods, and not the estimation accuracy of single-tree models, so that subsampling and honest splitting would induce otherwise irrelevant overhead.
Appendix D Simulation Figures for Section 6
VCM experiments.
@@ Appendix D Simulation Figures for Section 6
Figures 6789 illustrate the absolute fit times for the GRF-FPT algorithm under the four VCM settings of $\theta^{*}_{k}(x)$ described in Table 2. Across all settings and all dimensions, GRF-FPT is consistently several factors faster than GRF-grad. The speedup factor is summarized in Figure 4, which illustrates the relative speedup of GRF-FPT, calculated as the ratio of GRF-grad fit times over GRF-FPT fit times. Consistent with the observations in Section 5, we find that the speed advantage of GRF-FPT increases as the dimension of the target increases. Moreover, Figure 10 shows that this speed advantage comes at no penalty with respect to accuracy. Across all settings, GRF-FPT estimates $\hat{\theta}(x)$ display similar MSE results to estimates made by GRF-grad for $K=4$-dimensional targets $\theta^{*}(x)$ for VCMs. Details of the MSE estimation experiment are discussed in Section 6.
@@
Figure 6: Absolute fit times under Setting 1 for VCM timing experiments.
Figure 7: Absolute fit times under Setting 2 for VCM timing experiments.
Figure 8: Absolute fit times under Setting 3 for VCM timing experiments.
Figure 9: Absolute fit times under Setting 4 for VCM timing experiments.
@@
Figure 10: Estimates of MSE $\mathbb{E}[\|\theta^{*}(X)-\hat{\theta}(X)/K\|^{2}_{2}]$ for VCM for $K=4$ dimensional (top) and $K=256$ dimensional targets (bottom).
HTE experiments.
Figures 1314151617 illustrate the absolute fit times for the GRF-FPT algorithm under the five HTE settings of $\theta^{*}_{k}(x)$ and $\pi_{k}(x)$ described in Table 3. We find that GRF-FPT is consistently faster than GRF-grad. The speedup factor is summarized in Figure 18, which illustrates the relative speedup of GRF-FPT, calculated as the ratio of GRF-grad fit times over GRF-FPT fit times. As was seen for VCM experiments, the speed advantage of GRF-FPT scales with the dimensionality $K$ of the target. Moreover, Figure 19 shows that shows that this speed advantage comes at no penalty with respect to accuracy. Across all settings, for both $K=4$ and $K=256$-dimensional targets, the HTE estimates $\hat{\theta}(x)$ made by GRF-FPT display similar MSE to estimates made by GRF-grad. Details of the MSE estimation experiment are discussed in Section 6.
Figure 11: Absolute fit times for HTE timing experiments.
Figure 12: Speedup factor for HTE timing experiments.
Figure 13: Absolute fit times under Setting 1 for HTE timing experiments.
Figure 14: Absolute fit times under Setting for HTE timing experiments.
Figure 15: Absolute fit times under Setting 3 for HTE timing experiments.
Figure 16: Absolute fit times under Setting 4 for HTE timing experiments.
Figure 17: Absolute fit times under Setting 5 for HTE timing experiments.
@@
Figure 18: Speedup factor for HTE timing experiments. The settings for the HTE functions $\theta^{*}_{k}(\cdot)$ are found in Table 3.
@@
Figure 19: Estimates of the MSE $\mathbb{E}[\|\theta^{*}(X)-\hat{\theta}(X)/K\|^{2}_{2}]$ for HTE estimates $\hat{\theta}(x)$ for $K=4$ dimensional (top) and $K=256$ dimensional targets (bottom).
Appendix E Additional Examples of Pseudo-Outcomes
Nonparametric regression.

### v4

@@ C.1 Honest subsampling
$\alpha_{i}(x)\coloneqq\frac{1}{B}\sum^{B}_{b=1}\alpha_{bi}(x),\quad\text{for}\quad\alpha_{bi}(x)\coloneqq\frac{\mathds{1}(X_{i}\in L_{b}(x))}{|L_{b}(x)|},\qquad i=1,\ldots,n,$ (63)
@@
$\mathbb{E}[\alpha_{bi}(x)\mid O_{i},X_{i}]=\mathbb{E}[\alpha_{bi}(x)\mid X_{i}].$
@@
    $\mathbb{E}[\alpha_{bi}(x)\mid O_{i},X_{i}]=\mathbb{E}[\alpha_{bi}(x)\mid X_{i}]=0,\qquad\text{for all}\quad i\notin\mathcal{J}^{(b)}_{2}.$
@@
    $\mathbb{E}[\alpha_{bi}(x)\mid O_{i},X_{i}]=\mathbb{E}[\alpha_{bi}(x)\mid X_{i}],\qquad\text{for all}\quad i\in\mathcal{J}^{(b)}_{2}.$
@@ C.4 Simulation Details
Implementation details.
We implement the GRF-FPT algorithm in a fork of grf (Tibshirani et al., 2024) available at https://github.com/dfleis/grf. The functions grf::lm_forest and grf::multi_arm_causal_forest provide an easy to use interface for VCM and HTE estimation, respectively, and we allow the choice GRF-FPT1, GRF-FPT2, or GRF-grad to be controlled via the method argument. Code and data for reproducing all experiments and figures are available at https://github.com/dfleis/grf-experiments.
Data-generating settings.
The different setting for the target effects $\theta_{k}^{*}(x)$ include a sparse linear setting, a sparse logistic setting with interaction, a dense logistic setting, and a random function generator setting. Tables 2 and 3 provide the details of each regime for VCM and HTE experiments, respectively, for the data-generating model (28). These tables also summarize the different settings used to generate the $K$-dimensional regressors $W_{i}=(W_{i,1},\ldots,W_{i,K})^{\top}$. For VCM experiments, $W_{i,k}\sim\mathcal{N}(0,1)$ for all $k=1,\ldots,K$. For HTE experiments, $W_{i}\mid X_{i}=x\sim\text{Multinomial}(1,(\pi_{1}(x),\ldots,\pi_{K}(x)))$, where $\pi_{k}(x)$ denotes the underlying probability the sample is observed as having treatment level $k\in\{1,\ldots,K\}$.
@@
$K$ 4; 16;
$n$ 1000; 4000
$\dim(\mathcal{X})$ 2
nTrees 100; 500
Table 1: Parameter values for VCM and HTE experiments in Section 6. Target/regressor dimension $K$, number of observations $n$, dimension of the auxiliary variables $\dim(\mathcal{X})$, and number of trees nTrees. Experiments include a large-$n$ setting (left table) and a small-$n$ setting (right table).
@@ 4 $\theta^{*}_{k}(x)=\text{RFG}(x)$ $\mathcal{N}(0,1)$
Table 2: Settings for the true effects $\theta^{*}_{k}(\cdot)$ and the regressors $W_{i,k}$ for VCM experiments in Section 6. The function $\varsigma(u)\coloneqq 1+(1+e^{-20(u-1/3)})^{-1}$ is a logistic-type function in (Athey et al., 2019). The random function generator $\text{RFG}(x)$ is described in Appendix C.4.
@@
Table 3: Settings for the underlying treatment effects $\theta^{*}_{k}(\cdot)$ and treatment probabilities $\pi_{k}(x)$ for HTE experiments in Section 6. The function $\varsigma(u)\coloneqq 1+(1+e^{-20(u-1/3)})^{-1}$ is a logistic-type function used in (Athey et al., 2019). The random function generator $\text{RFG}(x)$ is described in Appendix C.4.
Random function generator.
@@
The effect functions $\theta^{*}_{k}(x)=\text{RFG}(x)$ under VCM Setting 4 (in Table 2) and HTE Setting 5 (in Table 3) follow the random function generator design of Friedman (2001). The idea is to measure the performance of the estimator under a variety of randomly generated targets. Each $\theta^{*}_{k}(\cdot)$ is randomly generated as a linear combination of functions $\{g_{\ell}(\cdot)\}_{\ell}^{20}$ of the form
$\theta^{*}_{k}(x)=\sum_{\ell=1}^{20}a_{\ell}g_{\ell}(z_{\ell}),$
@@
where the coefficients $\{a_{\ell}\}_{\ell=1}^{20}$ are randomly generated from a uniform distribution $a_{\ell}\sim\mathcal{U}([-1,1])$. Each $g_{l}(z_{l})$ is a function of a randomly selected $p_{\ell}$-size subset of the $p$-dimensional variable $x$, where the size of each subset $p_{\ell}$ is randomly chosen by $p_{\ell}=\min(\left\lfloor 1.5+r_{\ell}\right\rfloor,p)$, and $r_{\ell}$ is generated from an exponential distribution with mean $2$, $r_{\ell}\sim\mathrm{Exp}(0.5)$. Each $g(z_{\ell})$ uses a $p_{\ell}$-sized random subset $z_{\ell}\in\mathbb{R}^{p_{\ell}}$ of the $p$-dimensional input $x\in\mathbb{R}^{p}$:
$z_{\ell}\coloneqq\left(x_{\phi_{\ell}(1)},\ldots,x_{\phi_{\ell}(p_{\ell})}\right)\in\mathbb{R}^{p_{\ell}},$
such that $\{\phi_{\ell}(1),\ldots,\phi_{\ell}(p_{\ell})\}$ is a length-$p_{\ell}$ permutation of indices drawn from $\{1,\ldots,p\}$, without replacement. The functions $g_{\ell}(\cdot)$ are Gaussian functions of the $p_{\ell}$ sampled variables:
$g_{\ell}(z_{\ell})\coloneqq\exp\left\{-\frac{1}{2}(z_{\ell}-\mu_{\ell})^{\top}\mathbf{V}_{\ell}(z_{\ell}-\mu_{\ell})\right\},$
where the mean vector $\mu_{\ell}\in\mathbb{R}^{p_{\ell}}$ is randomly generated from a standard multivariate Gaussian, $\mu_{\ell}\sim\mathcal{N}_{p_{\ell}}({\bf 0},\mathbb{I})$. The $p_{l}\times p_{l}$ covariance matrix $\mathbf{V}_{l}$ are formed through the spectral decomposition:
$\mathbf{V}_{\ell}=\mathbf{U}_{\ell}\mathbf{D}_{\ell}\mathbf{U}_{\ell}^{\top},$
@@
where $\mathbf{U}_{\ell}$ is a random $p_{\ell}\times p_{\ell}$ orthonormal matrix and $\mathbf{D}_{\ell}\coloneqq\text{diag}(d_{1,\ell},\ldots,d_{p_{\ell},\ell})$ with diagonal entries $d_{j,\ell}$ generated from a uniform distribution according to $\sqrt{d_{j,\ell}}\sim\mathcal{U}(0.1,2.0)$.
Appendix D Additional Simulations
D.1 Settings for the criterion value experiment in Section 3.4
The criterion value experiment in Section 3.4 was run under a varying coefficient model of the form
$Y_{i}\coloneqq W_{i}^{\top}\theta^{*}(X_{i})+\epsilon_{i},\qquad\epsilon_{i}\sim\mathcal{N}(0,0.5^{2}),$ (64)
where the regressors $W_{i}$ were generated as bivariate standard Gaussian samples $W_{i}\sim\mathcal{N}_{2}({\bf 0},\mathbb{I})$ and the auxiliary covariates were generated as standard uniform samples $X_{i}\sim\mathcal{U}(0,1)$. The data-generating coefficient functions were $\theta^{*}(x)\coloneqq(\sin(2\pi x),\;x)$ and the criterion values were computed based on $n=1000$ samples following (64).
D.2 Supporting experiments for Section 6
Multicollinearity in auxiliary covariates.
We conducted a VCM experiment with highly correlated auxiliary covariate features. We ran a modified version of VCM Setting 3 by generating auxiliary covariates as $X_{i}\sim\mathcal{N}({\bf 0},\Sigma)$, where $[\Sigma]_{j,k}=\omega^{|j-k|}$ for $\omega\in\{0,0.5,0.9\}$. Table 4 provides a clear summary of the computational performance of GRF-FPT relative to GRF-grad and statistical accuracy (MSE – multiplied by 100 for readability). All experiments were run over a forest of 10 trees and MSE estimates were computed over 50 replications of the model and evaluated on a separate set of $n=5,000$ samples, carried out using GRF-FPT2 and GRF-grad. These results demonstrate clearly that GRF-FPT remains robust, stable, and computationally efficient, even under high multicollinearity in $X_{i}$.
$\dim(\mathcal{X})$ $n$ $K$ $\omega$ Speedup $100\times\text{MSE}$ grad $100\times\text{MSE}$ FPT2
5 10,000 64 0.00 2.55 16.60 16.83
5 10,000 64 0.50 2.53 15.48 15.47
5 10,000 64 0.90 2.35 10.95 11.09
Table 4: Effect of multicollinearity in the auxiliary covariates $X_{i}$ on the relative computational gain of GRF-FPT2, as well as the statistical accuracy of both GRF-FPT and GRF-grad estimators.
Subsampling ratio.
We carried out an experiment to show that the subsample proportion does not affect the computational advantage or statistical accuracy of GRF-FPT relative to GRF-grad. We varied the subsampling ratio $s/n\in\{0.25,0.50,0.75\}$ under VCM Setting 3 over a forest of 10 trees carried out using GRF-FPT2 and GRF-grad. Table 5 summarizes our results, averaged over 50 replications of the model, with a test set of 5,000 samples. These results show clearly that the statistical accuracy (MSE) of GRF-FPT2 relative to GRF-grad does not depend strongly on the subsample ratio.
$\dim(\mathcal{X})$ $n$ $K$ $s/n$ Speedup $100\times\text{MSE}$ grad $100\times\text{MSE}$ FPT2
2 10,000 64 0.25 2.77 2.86 2.90
2 10,000 64 0.50 3.10 2.91 2.90
2 10,000 64 0.75 2.98 3.21 3.19
Table 5: Effect of the subsampling ratio $s/n$ on the relative computational gain of GRF-FPT2, as well as the statistical accuracy of both GRF-FPT and GRF-grad estimators.
Large sample size.
We ran additional experiments to clearly show how our method scales for very large datasets. Using a forest of 10 trees, we tested our method on VCM Setting 3 with sample sizes up to $n=500,000$, carried out using GRF-FPT2 and GRF-grad. The results are summarized in Table 6 and demonstrate that, even as the dataset grows very large, our method consistently remains faster than GRF-grad. While the relative speedup slightly decreases at first, it stabilizes towards a consistent advantage as grows $n$ grows sufficiently large, suggesting that the advantage is not bottlenecked by $n$ and maintains a robust advantage at scale.
$\dim(\mathcal{X})$ $K$ $n$ Speedup
2 256 10,000 4.54
2 256 20,000 3.59
2 256 50,000 3.49
2 256 100,000 3.11
2 256 200,000 3.04
2 256 500,000 3.08
Table 6: Effect of increasing sample sizes $n$ on the relative computational gain of GRF-FPT2.
D.3 Supporting figures for Section 6
D.3.1 VCM experiments
Large $n$ VCM simulations.
@@ D.3.1 VCM experiments
Figure 5 illustrates the absolute fit times for the GRF-FPT algorithms under the four VCM settings for $\theta^{*}_{k}(x)$ described in Table 2 over the large-$n$ settings in Table 1. Across all settings and all dimensions, GRF-FPT is consistently several factors faster than GRF-grad. The speedup factor is summarized in Figure 3, which illustrates the relative speedup of GRF-FPT, calculated as the ratio of GRF-grad fit times over GRF-FPT fit times. Consistent with the observations in Section 5, we find that the speed advantage of GRF-FPT increases as the dimension of the target increases.
Figure 6 shows that this speed advantage comes while performing comparably to GRF-grad in terms of statistical accuracy. Across all settings for VCMs with $K=4$ dimensional targets, the MSE estimates from GRF-FPT is highly similar to the MSE estimates of GRF-grad, while for $K=256$ dimensional targets one sees more variation in MSE estimates across the methods. This effect likely reflects the increased variance associated with high-dimensional estimation. In some cases we see GRF-FPT1 slightly outperform both GRF-FPT2 and GRF-grad, in other cases we see GRF-grad slightly outperform both GRF-FPT methods, and in others GRF-FPT2 yields the lowest MSE. One sees that these differences are typically small. The key benefit we emphasize is that GRF-FPT is able to achieve nearly identical statistical accuracy with a substantial improvement in computational speed.
Small $n$ VCM simulations.
Figures 8 and 7 illustrate the absolute fit times and relative speed advantage, respectively, of GRF-FPT under the VCM design of $\theta_{k}^{*}(x)$ over the small-$n$ settings. One sees that even when $n$ is more modest, GRF-FPT consistently offers a computational advantage over GRF-grad, with possible outliers under VCM Setting 2 at $K=4$. We believe this negative relative advantage to be caused by random fluctuations in computation and are not representative of the FPT algorithm itself, particularly in light of the fact that the negative effect vanishes when the number of trees increases from 100 to 500. As one would expect based on the large-$n$ results, the relative advantage tends to increase with increasing $K$, and generally stabilizes with increasing $n$. Figure 9 shows that the GRF-FPT speed advantage does not come at any material cost in statistical accuracy, with similar performance to GRF-grad across all settings.
@@
Figure 5: Absolute fit times for VCM timing experiments under the settings in Table 2 and large-$n$ settings in Table 1.
@@
Figure 6: Estimates of MSE $\mathbb{E}[\|\theta^{*}(X)-\hat{\theta}(X)/K\|^{2}_{2}]$ for VCM for $K=256$ dimensional (top) and $K=4$ dimensional targets (bottom) under the large-$n$ settings in Table 1.
@@
Figure 7: Speedup factor for GRF-FPT in comparison to GRF-grad for VCM timing experiments under the small-$n$ settings in Table 1.
Figure 8: Absolute fit times for VCM timing experiments under the settings in Table 2 and small-$n$ settings in Table 1.
@@
Figure 9: Estimates of MSE $\mathbb{E}[\|\theta^{*}(X)-\hat{\theta}(X)/K\|^{2}_{2}]$ for VCM for $K=16$ dimensional (top) and $K=4$ dimensional targets (bottom) under the small-$n$ settings in Table 1.
D.4 HTE experiments
Large $n$ HTE simulations.
Figure 11 illustrates the absolute fit times for the GRF-FPT algorithm under the five HTE settings of $\theta^{*}_{k}(x)$ and $\pi_{k}(x)$ described in Table 3 over the large-$n$ settings in Table 1. We find that GRF-FPT is consistently faster than GRF-grad. The speedup factor of GRF-FPT relative to GRF-grad is summarized in Figure 10, calculated as the ratio of GRF-grad fit times over GRF-FPT fit times. As was seen for VCM experiments, the speed advantage of GRF-FPT scales with the dimensionality $K$ of the target. One sees from both Figures 10 and 11 that GRF-FPT’s computational advantage is less dramatic than under the VCM experiments. This can be understood based on the fact that the VCM regressors $W_{i}$ are continuous while the HTE regressors represent binary indicators. Continuous regressors provide more granularity when fitting the child statistics $\tilde{\theta}_{C_{j}}$, and as a result provide a larger set of candidate splits over the covariates. Nevertheless, one sees in Figure 10 that the FPT splitting mechanism is still up to 1.5$\times$ faster under the largest regressor setting $K=256$, with a more modest, but persistent savings across all settings.
The statistical benchmarks for our HTE experiments are shown in Figure 12. Consistent with the VCM experiments, one sees that the computational advantage of GRF-FPT does not come at the cost of in terms of its statistical accuracy.
Small $n$ HTE simulations.
Figures 13 and 14 summarize the relative speed advantage and absolute fit times for the GRF-FPT algorithm under the small-$n$ HTE design. Consistent with the large-$n$ HTE experiments the FPT2 mechanism sees a stable computational advantage across all settings, with an increasing effect in increasing $K$, while the FPT1 mechanism displays a persistent advantage for $K=16$ and comparable computational performance for $K=4$. The more modest relative advantage for the small-$n$ experiments is itself consistent with the VCM small-$n$ experiments, owing in large part due to the smaller values of $K$. Figure 15 compares the statistical performance of GRF-FPT to GRF-grad, with no material difference between either GRF-FPT1, GRF-FPT2, or GRF-grad’s estimation accuracy.
Figure 10: Speedup factor for GRF-FPT in comparison to GRF-grad for HTE timing experiments under the large-$n$ setting in Table 1.
Figure 11: Absolute fit times for HTE timing experiments under the settings in Table 3 and large-$n$ settings in Table 1.
Figure 12: Estimates of MSE $\mathbb{E}[\|\theta^{*}(X)-\hat{\theta}(X)/K\|^{2}_{2}]$ for HTE for $K=256$ dimensional (top) and $K=4$ dimensional targets (bottom) under the large-$n$ settings in Table 1.
Figure 13: Speedup factor for GRF-FPT in comparison to GRF-grad for HTE timing experiments under the small-$n$ settings in Table 1.
Figure 14: Absolute fit times for HTE timing experiments under the settings in Table 3 and small-$n$ settings in Table 1.
Figure 15: Estimates of MSE $\mathbb{E}[\|\theta^{*}(X)-\hat{\theta}(X)/K\|^{2}_{2}]$ for HTE for $K=16$ dimensional (top) and $K=4$ dimensional targets (bottom) under the small-$n$ settings in Table 1.
Appendix E Additional Examples
E.1 Pseudo-outcomes for nonparametric regression